# When Data is the Algorithm: A Systematic Study and Curation of Preference Optimization Datasets

### v1

@@ A.1 General Post-Training Methods
The goal of preference fine-tuning is is to align the SFT model’s output distribution with behavior that appeals to human preferences or adheres to task-specific objectives. This is commonly achieved by leveraging a learned reward signal or with explicit preference annotations to steer the model toward generating responses that reflect pre-defined behavior such as increased helpfulness, harmlessness, honesty, and style. Widely used methods for preference-based alignment include Proximal Policy Optimization (PPO) [schulman2017proximalpolicyoptimizationalgorithms] and Direct Preference Optimization (DPO) [rafailov2024directpreferenceoptimizationlanguage].
@@
Recent work has explored RL methods that enhance a model’s capacity for structured reasoning. For instance, Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] has been proposed to elicit deeper, multi-step thinking by refining reward signals through group-based preference aggregation. In parallel, several reasoning-centric datasets have been introduced [bercovich2025llamanemotronefficientreasoningmodels, sky_t1_2025, bespoke_stratos, slam-distillation-from-r1, still], focusing on task formats that explicitly promote step-by-step problem solving and reflective thought processes.
@@ A.2 Focus on DPO
The primary goal of this paper is to analyze the quality and composition of preference optimization datasets while keeping the overall training procedure fixed, enabling a fair comparison. In particular, we focus on DPO, which has become the most widely adopted approach for preference-based alignment due to its simplicity and efficiency [ivison2024unpackingdpoppodisentangling]. Unlike PPO that requires both reward modeling and costly policy rollouts, DPO operates directly on preference pairs, making it well suited for dataset-centric comparisons given the large availability of open-source DPO corpora.
Similar efforts have been pursued by tulutalk for post-training with SFT, demonstrating that task-aware dataset design can improve performance while reducing overall dataset size. However, the diversity of existing post-training methods, including SFT, PPO, and others, implies that dataset requirements are highly method-dependent, complicating direct comparisons across approaches. For instance, unlike SFT, DPO is based on the Bradley-Terry formulation, which relies on pairwise preference samples for optimization. As a result, curation strategies developed for SFT may not be directly applicable to DPO datasets. To address this gap, we focus specifically on post-training with DPO, analyzing the role of preference rewards and assessing the reliability of pre-annotated scores, thus going beyond binary preference labels to enable more insightful dataset construction.
@@
Appendix B DPO Datasets
@@
TuluDPO. lambert2025tulu3pushingfrontiers originally curated TuluDPO for post-training Llama-3.1 models [grattafiori2024llama3]. The dataset consists of 272,898 preference pairs drawn from several distinct sources, consolidating on- and off-policy preference data from over seven constituent datasets, including UltraFeedback [cui2024ultrafeedbackboostinglanguagemodels] and WildChat [zhao2024wildchat]. The corresponding responses were generated using a diverse set of open and frontier LLMs. While TuluDPO represents one of the most comprehensive, general-purpose DPO corpora available, its construction inherits samples from several upstream mixtures without per-sample metadata, making fine-grained analysis and targeted reuse difficult without additional annotations.
@@
ORPO. orpo curated ORPO as a mixture of open-source preference pairs designed for training with Optimized Rejection-based Preference Optimization (ORPO) [orpo_paper] and DPO. It aggregates 44,218 high-scoring preference pairs from several high-quality sources, particularly from Argilla [argilla_hub], with chosen responses filtered to meet minimum reward thresholds. ORPO offers a broad distribution of general, factual, and safety-aligned instructions for training general-purpose preference models.
@@
UltraFeedback. cui2024ultrafeedbackboostinglanguagemodels released curated preference pairs for over 60,000 prompts, spanning instruction following, truthfulness, honesty, and helpfulness. Instructions were sourced from high-quality public datasets such as FLAN [longpre2023flan], Evol-Instruct [luo2024wizardcoder], and TruthfulQA [lin-etal-2022-truthfulqa], and their completions were sampled from 17 diverse LLMs, including GPT-4 [openai2024gpt4ocard]. While UltraFeedback does provide scalar preference scores (assessed via GPT-4), they remain inconsistent, for example, assigning the same reward score for both chosen and discarded responses. This can occur when multiple responses demonstrate similar performance and the preference pair is thus selected arbitrarily.
@@
HelpSteer. helpsteer introduced HelpSteer2 (referred to as just HelpSteer) as a high-quality instruction following dataset with human annotations for helpfulness, correctness, coherence, complexity, and verbosity. The small-scale dataset comprises 10,681 prompts with responses collected primarily from ShareGPT. Each response is scored by over 1,000 human annotators on a fine-grained Likert-scale, thus offering explicit, multi-dimensional preference signals rather than binary choices. These annotations can be used to construct preference pairs, for example, based on mean scores, as done in our work.
@@
Code-Preference-Pairs. codepreferences introduced this synthetic code dataset to train models for bug identification and correction. Each sample contains two code variants: an accepted version with correct functionality and inline explanations, and a discarded version where bugs are surgically introduced without explicit indicators. The dataset spans over 55,000 preference pairs across multiple languages and coding problems, with samples derived from several sources, including Evol-Instruct [luo2024wizardcoder].
@@
We specifically selected these datasets due to their prominence and frequent appearance in open-source community blogs and dataset hubs [mlabonne_llm-datasets_2025]. While we acknowledge that new, potentially better-performing datasets are continuously being released, our analysis, annotation pipeline, and data curation recipes are generalizable and can, in principle, be applied to any DPO dataset.
@@ C.1 General Overview
Magpie [xu2024magpie] is a self-synthesis framework that derives alignment-oriented annotations from open-weight, instruction-tuned language models without requiring human labels or handcrafted seed prompts. Although the framework is also capable of generating new instruction-response pairs, in this work we rely exclusively on its annotation capabilities.
@@
A key feature of Magpie is its modularity: any instruction-tuned LLM can, in principle, serve as the judge model. In its default configuration, Magpie relies on Llama-3-8B-Instruct [grattafiori2024llama3] for general annotation and Llama-Guard 2 [metallamaguard2] for safety assessment. In our pipeline, we use Llama-3.3-70B-Instruct [grattafiori2024llama3] as the judge model given its demonstrated reliability, particularly for quality and difficulty annotations, as assessed by tulutalk. In addition, we also incorporate the error-tolerant JSON parsing extensions to Magpie from tulutalk, which help handle formatting inconsistencies and free-form outputs during the annotation process.
@@ C.2 Preference Reward Assessment
We generate preference rewards by adapting Magpie to evaluate the response quality of each completion. To this end, we rely on an independent, specialized reward model that has been fine-tuned on multiple preference datasets. Specifically, we use FsfairX, a Llama-3-8B-Instruct-based reward model introduced by reward_model, which was trained using the Bradley-Terry formulation on a diverse set of high-quality preference datasets, including HH-RLHF [hhrlhf], SHP [shp], and UltraFeedback cui2024ultrafeedbackboostinglanguagemodels. This reward model has demonstrated strong alignment with human preference judgments, making it a reliable signal for quantifying how well a model completion satisfies the corresponding instruction.
@@
While we acknowledge that other reward models can be used as alternatives, such as UltraRM-13B [cui2024ultrafeedbackboostinglanguagemodels], Snorkel-Mistral-PairRM-DPO [snorkel_ai], ArmoRM [armoRM], Skywork-Reward [liu2024skywork], or more generally, pairwise preference models trained on high-quality preference datasets, reward_model demonstrated through various experiments that their Llama-based FsfairX model performs comparably or more reliably than these alternatives and aligns closely with human judgments, thereby justifying our choice. Nevertheless, we provide a side-by-side comparison to the Skywork reward model for TuluDPO and ORPO datasets in Fig. 6(a) and Fig. 6(b), respectively, showing overall similar distributions for the difference in preference rewards.
In particular, the similar shape of the distributions indicates that preference reward assessment follows the same general trend and does not fundamentally alter the evaluation outcome. At the same time, we observe that the FsfairX model produces fewer negative outliers and more stable margins, making it particularly well-suited for providing reliable training signals without over-amplifying differences. In contrast, the Skywork reward model exhibits both stronger positive separations and more negative outliers, suggesting higher variance. For our purposes, FsfairX provides the consistency and robustness required for a balanced assessment of preference rewards, with its margins further validated against human feedback in several experiments [reward_model].

### v2

@@ A.1 General Post-Training Methods
The goal of preference fine-tuning is is to align the SFT model’s output distribution with behavior that appeals to human preferences or adheres to task-specific objectives. This is commonly achieved by leveraging a learned reward signal or with explicit preference annotations to steer the model toward generating responses that reflect pre-defined behavior such as increased helpfulness, harmlessness, honesty, and style. Widely used methods for preference-based alignment include Proximal Policy Optimization (PPO) [Schulman et al., 2017] and Direct Preference Optimization (DPO) [Rafailov et al., 2023].
@@
Recent work has explored RL methods that enhance a model’s capacity for structured reasoning. For instance, Group Relative Policy Optimization (GRPO) [Shao et al., 2024] has been proposed to elicit deeper, multi-step thinking by refining reward signals through group-based preference aggregation. In parallel, several reasoning-centric datasets have been introduced [Bercovich et al., 2025, Team, 2025, Bespoke-Labs, 2025, Madhusudhan et al., 2025, Min et al., 2024], focusing on task formats that explicitly promote step-by-step problem solving and reflective thought processes.
@@ A.2 Focus on DPO
The primary goal of this paper is to analyze the quality and composition of preference optimization datasets while keeping the overall training procedure fixed, enabling a fair comparison. In particular, we focus on DPO, which has become the most widely adopted approach for preference-based alignment due to its simplicity and efficiency [Ivison et al., 2024]. Unlike PPO that requires both reward modeling and costly policy rollouts, DPO operates directly on preference pairs, making it well suited for dataset-centric comparisons given the large availability of open-source DPO corpora.
Similar efforts have been pursued by Djuhera et al. [2025] for post-training with SFT, demonstrating that task-aware dataset design can improve performance while reducing overall dataset size. However, the diversity of existing post-training methods, including SFT, PPO, and others, implies that dataset requirements are highly method-dependent, complicating direct comparisons across approaches. For instance, unlike SFT, DPO is based on the Bradley-Terry formulation, which relies on pairwise preference samples for optimization. As a result, curation strategies developed for SFT may not be directly applicable to DPO datasets. To address this gap, we focus specifically on post-training with DPO, analyzing the role of preference rewards and assessing the reliability of pre-annotated scores, thus going beyond binary preference labels to enable more insightful dataset construction.
@@
A.3 Theoretical Perspective on Data Quality
DPO can be viewed as optimizing a KL–regularized reward objective where the only supervision comes from pairwise preferences Rafailov et al. [2023]. Formally, given a reference policy $\pi_{\mathrm{ref}}$ and a dataset of tuples $(x,y_{c},y_{r})$ with preferred completion $y_{c}$ and rejected completion $y_{r}$, the DPO loss for a policy $\pi_{\theta}$ can be written as
$\mathcal{L}_{\mathrm{DPO}}(\theta)~=~\mathbb{E}_{(x,y_{c},y_{r})\sim\mathcal{D}}\bigl[-\log\sigma\bigl(\beta\,m_{\theta}(x)\bigr)\bigr],\ m_{\theta}(x)~=~\log\frac{\pi_{\theta}(y_{c}\mid x)}{\pi_{\mathrm{ref}}(y_{c}\mid x)}-\log\frac{\pi_{\theta}(y_{r}\mid x)}{\pi_{\mathrm{ref}}(y_{r}\mid x)}\ ,$ (1)
where $\sigma$ is the logistic function and $\beta>0$ controls the KL strength. The quantity $m_{\theta}(x)$ is the log-likelihood margin between chosen and rejected responses under the implicit reward model
$r_{\theta}(x,y)~=~\beta\,\bigl(\log\pi_{\theta}(y\mid x)-\log\pi_{\mathrm{ref}}(y\mid x)\bigr)\ ,$ (2)
which DPO fits directly from preferences. Under the Bradley–Terry model, this implicit reward is provably equivalent (up to an additive baseline) to the reward learned in RLHF, and the optimal policy of DPO matches the KL–constrained RL optimum.
From this perspective, the data enter the optimization only through the distribution of margins $m_{\theta}(x)$ induced by the preference pairs. If preferences are consistent and the chosen response is substantially better than the rejected one, the margin is positive and large in magnitude, yielding a high signal-to-noise ratio (SNR) for the gradient. Conversely, incoherent or near-tied pairs induce small or even negative margins, which (i) reduce the effective SNR and (ii) generate conflicting gradients that pull $\pi_{\theta}$ in incompatible directions. Our empirical findings that 20–30% of pairs in popular datasets violate reward coherence, and that prompt quality and reward margins are strongly correlated, provide direct evidence for this mechanism.
A complementary view is given by the $\Psi$-preference optimization ($\Psi$-PO) framework of Azar et al. [2023], which treats learning from preferences as an offline contextual bandit problem. They define a general objective
$J_{\Psi}(\pi)~=~\mathbb{E}_{x\sim\rho}\mathbb{E}_{y\sim\pi(\cdot\mid x),\,y^{\prime}\sim\mu(\cdot\mid x)}\bigl[\Psi\bigl(p^{*}(y\succ y^{\prime}\mid x)\bigr)\bigr]~-~\tau\,D_{\mathrm{KL}}\!\bigl(\pi\,\|\,\pi_{\mathrm{ref}}\bigr)\ ,$ (3)
where $p^{*}(y\succ y^{\prime}\mid x)$ is the human preference probability, $\mu$ is a behavior policy, $\Psi$ is any non-decreasing function, and $\tau>0$ controls regularization. Under the Bradley–Terry model and the specific choice $\Psi(q)=\log\frac{q}{1-q}$, they show that $\Psi$-PO recovers both RLHF and DPO as special cases.
In this formulation, dataset quality is encoded directly in the preference probabilities. If prompts are underspecified, many pairs have $p^{*}(y_{c}\succ y_{r}\mid x)\approx 0.5$, so $\Psi\bigl(p^{*}(\cdot)\bigr)$, degrading the effective gradient. If preference labels are inconsistent (e.g., $p^{*}(y_{c}\succ y_{r}\mid x)<0.5$), updates push the policy toward worse responses. If entire task regions are underrepresented, the expectation over $x\sim\rho$ underweights those regimes, regardless of how clean individual pairs are.
Our Magpie-based annotations and curation recipe improve the $\Psi$-PO objective in three ways. First, higher prompt quality sharpens $p^{*}(y_{c}\succ y_{r}\mid x)$ by reducing ambiguity, increasing the expected margin $m_{\theta}(x)$ and lowering its variance—consistent with our observation that better inputs yield larger reward gaps. Second, reward coherence removes pairs where the rejected completion scores higher ($r(x,y_{r})\geq r(x,y_{c})$), eliminating preference inconsistencies that can cause overfitting and instability in margin-based objectives like DPO [Azar et al., 2023]. This reduces systematic label noise and stabilizes the margin distribution. Third, task-aware curation reshapes the context distribution $\rho$ by rebalancing task categories. Since instruction-following drives cross-task gains while math and code depend more on absolute scale, this effectively reweights training toward contexts with abundant, high-quality, high-margin preferences that best support generalization.
In summary, existing theory already provides a principled link between data quality and DPO-style objectives: (i) DPO implicitly fits a reward model whose training signal is governed by the log-likelihood margin $m_{\theta}(x)$, and (ii) $\Psi$-PO shows that the shape of the preference probabilities $p^{*}$, as influenced by prompts, labels, and task mix, controls the effective optimization landscape. Our empirical work complements this by (a) measuring prompt quality and reward coherence at scale, (b) showing that misaligned and low-margin pairs are common and harmful, and (c) demonstrating that principled filtering on these signals yields consistent improvements across models and benchmarks.
Appendix B Prior and Related Work
B.1 DPO Datasets
@@
TuluDPO. Lambert et al. [2025] originally curated TuluDPO for post-training Llama-3.1 models [Grattafiori et al., 2024]. The dataset consists of 272,898 preference pairs drawn from several distinct sources, consolidating on- and off-policy preference data from over seven constituent datasets, including UltraFeedback [Cui et al., 2024] and WildChat [Zhao et al., 2024]. The corresponding responses were generated using a diverse set of open and frontier LLMs. While TuluDPO represents one of the most comprehensive, general-purpose DPO corpora available, its construction inherits samples from several upstream mixtures without per-sample metadata, making fine-grained analysis and targeted reuse difficult without additional annotations.
@@
ORPO. Labonne [2024] curated ORPO as a mixture of open-source preference pairs designed for training with Optimized Rejection-based Preference Optimization (ORPO) [Hong et al., 2024] and DPO. It aggregates 44,218 high-scoring preference pairs from several high-quality sources, particularly from Argilla [Argilla, 2025], with chosen responses filtered to meet minimum reward thresholds. ORPO offers a broad distribution of general, factual, and safety-aligned instructions for training general-purpose preference models.
@@
UltraFeedback. Cui et al. [2024] released curated preference pairs for over 60,000 prompts, spanning instruction following, truthfulness, honesty, and helpfulness. Instructions were sourced from high-quality public datasets such as FLAN [Longpre et al., 2023], Evol-Instruct [Luo et al., 2024], and TruthfulQA [Lin et al., 2022], and their completions were sampled from 17 diverse LLMs, including GPT-4 [Hurst et al., 2024]. While UltraFeedback does provide scalar preference scores (assessed via GPT-4), they remain inconsistent, for example, assigning the same reward score for both chosen and discarded responses. This can occur when multiple responses demonstrate similar performance and the preference pair is thus selected arbitrarily.
@@
HelpSteer. Wang et al. [2024c] introduced HelpSteer2 (referred to as just HelpSteer) as a high-quality instruction following dataset with human annotations for helpfulness, correctness, coherence, complexity, and verbosity. The small-scale dataset comprises 10,681 prompts with responses collected primarily from ShareGPT. Each response is scored by over 1,000 human annotators on a fine-grained Likert-scale, thus offering explicit, multi-dimensional preference signals rather than binary choices. These annotations can be used to construct preference pairs, for example, based on mean scores, as done in our work.
@@
Code-Preference-Pairs. Vezora [2024] introduced this synthetic code dataset to train models for bug identification and correction. Each sample contains two code variants: an accepted version with correct functionality and inline explanations, and a discarded version where bugs are surgically introduced without explicit indicators. The dataset spans over 55,000 preference pairs across multiple languages and coding problems, with samples derived from several sources, including Evol-Instruct [Luo et al., 2024].
@@
We specifically selected these datasets due to their prominence and frequent appearance in open-source community blogs and dataset hubs [Labonne, 2025]. While we acknowledge that new, potentially better-performing datasets are continuously being released, our analysis, annotation pipeline, and data curation recipes are generalizable and can, in principle, be applied to any DPO dataset.
B.2 Comparison with Related Work
Quantifying data quality is fundamentally non-trivial and prior works differ from ours in how they define what constitutes a “high-quality” preference dataset. We highlight these distinctions and provide new experimental evidence showing that UltraMix’s non-trivial combination of complementary signals, rather than quality alone, is essential for achieving the best performance.
Unpacking DPO/PPO.
[Ivison et al., 2024] compares datasets only monolithically and draws two conclusions: (a) synthetic datasets outperform human-annotated ones, and (b) datasets created using multi-aspect annotations outperform those with aggregated labels. However, their analysis does not examine per-sample quality, nor does it consider post-hoc curation. Our work goes beyond such monolithic comparisons by conducting a per-sample diagnosis, i.e., we look inside datasets and identify which concrete samples are subpar. We then leverage these sample-level annotations to propose a principled curation recipe. To this end, we provide the first large-scale quantification of prompt quality and reward coherence across widely used DPO datasets. Our results eventually show that even “high-quality” datasets (e.g., TuluDPO, UltraFeedback) contain 20–30% of samples with contradicting preference order.
AIR.
[He et al., 2025] focuses on creating a new dataset from scratch (AIR), whereas our work targets post-hoc curation, which is a crucial distinction. In particular, AIR extends the UltraFeedback pipeline to construct new datasets based on moderate reward margins and high absolute scores. In contrast, our workflow, which combines detailed annotations with a principled curation recipe, is designed specifically for post-hoc optimization. Consequently, AIR cannot be applied retroactively to existing datasets without regenerating and reannotating all samples, whereas UltraMix can. AIR further measures prompt quality by (1) generating multiple responses per prompt using diverse LLMs, (2) assigning LLM-based numerical scores to each response, and (3) defining prompt quality via the variance of these scores. This procedure is feasible only when constructing datasets from scratch and incurs substantial computational cost due to the need for many LLMs. By contrast, our post-hoc curation framework quantifies prompt quality through systematic annotations (e.g., clarity, reward coherence) without multi-LLM generation, making UltraMix both scalable and efficient.
RIP.
[Yu et al., 2025] investigates ways to filter preference data by maximizing reward differences. However, our approach instead evaluates reward coherence, that is, whether the chosen answer is actually better than the rejected one. Specifically, we show that relying on reward signals alone is insufficient. Our recipe thus balances prompt-level, reward-level, and task-level filtering. To further strengthen this point, we provide a direct comparison to RIP-style filtering in Table 5. We report results for RIP-only (filtering by 75th-percentile reward gap), RIP+UltraMix-Quality-Filter, and the full UltraMix recipe. The results show that RIP-only underperforms even when UltraMix’s quality filter is applied on top, which suggests that many pairs have similar reward values. Thus, discarding them solely based on reward gaps harms performance. Consequently, preference order is the more informative signal and our combined recipe yields a superior data filter.
Table 5: Comparison of RIP-style filtering versus UltraMix variants for Llama-3.1-8B-TuluSFT and Qwen-2.5-7B-TuluSFT. Benchmarks are grouped by category. Best scores per row are in bold.
Llama-3.1-8B-TuluSFT Qwen-2.5-7B-TuluSFT
Benchmark RIP-only RIP+UM-QF UltraMix RIP-only RIP+UM-QF UltraMix
Knowledge
MMLU (5-shot) 62.93 63.27 64.61 73.10 73.53 74.01
TruthfulQA (0-shot) 61.69 61.45 63.32 54.47 57.60 58.00
ARC-C (25-shot) 57.83 57.63 58.70 60.33 62.12 62.43
GSM8K (5-shot) 77.27 79.98 82.48 76.85 79.19 82.70
Commonsense
HellaSwag (10-shot) 63.35 63.43 64.82 62.67 63.59 63.89
WinoGrande (5-shot) 75.17 74.98 77.06 72.98 73.69 74.64
Reasoning
MMLU-Pro (5-shot) 28.99 28.50 30.96 43.48 43.16 44.65
BBH (3-shot) 43.93 44.68 44.96 48.08 50.96 51.76
GPQA (0-shot) 30.12 29.94 31.87 31.28 31.63 33.03
MuSR (0-shot) 42.42 42.43 44.02 47.19 47.87 48.82
Instruction Following
IF-Eval (0-shot) 79.31 79.38 81.13 78.03 77.28 79.88
Math
MATH (4-shot) 21.53 21.22 23.56 43.13 47.55 49.55
Code (pass@1)
HumanEval 65.71 65.61 69.05 80.49 78.05 82.27
HumanEval+ 45.77 45.76 48.08 61.83 60.49 63.05
Leaderboards
Open LLM Leaderboard 1 66.37 66.79 68.50 66.73 68.29 69.28
Open LLM Leaderboard 2 41.05 41.02 42.75 48.53 49.74 51.28
Overall 54.00 54.16 56.04 59.56 60.48 62.05
To further eliminate potential data bias, we provide additional results on the newer HelpSteer3 dataset [Wang et al., 2025]. To this end, we annotate the single-turn HelpSteer3 variant [^7] Annotated HelpSteer3 dataset: huggingface.co/datasets/aladinDJ/helpsteer3-single-turn-DPO-annotated using our DPO-adapted Magpie pipeline and conduct a comparison with RIP-style filtering on both Llama-3.1-8B and Qwen-2.5-7B. Specifically, we compare three configurations: HelpSteer3 (Baseline), HelpSteer3 (RIP), where we keep only samples with high reward gaps > 60th percentile, and HelpSteer3 (Pref), where we keep only samples where chosen reward > rejected reward.
We report the averaged scores across all leaderboard benchmarks (LB1 and LB2) in Table 6. The results show that preference coherence is the superior filtering signal, yielding better results compared to RIP-style filtering, which actually hurts performance (-1.08% on Llama and -0.67% on Qwen). These findings thus confirm for a separate dataset that high reward gaps do not guarantee high-quality training signals if the underlying preference direction is noisy or if the data distribution is skewed.
Table 6: DPO performance comparison of Llama and Qwen models on HelpSteer3 variants. Training follows the hyperparameter setup of Wang et al. [2025]. We report LB1, LB2, average score (Avg.), and the relative performance change ($\Delta$). Best scores within each model block are shown in bold.
Llama-3.1-8B-TuluSFT Qwen-2.5-7B-TuluSFT
Dataset LB1 LB2 Avg $\Delta$ LB1 LB2 Avg $\Delta$
HelpSteer3 (Base) 63.46 39.21 51.53 65.74 47.44 58.46
HelpSteer3 (RIP) 63.22 37.70 50.45 -1.08 65.42 46.59 57.79 -0.67
HelpSteer3 (Pref) 63.98 39.61 52.25 +0.72 66.02 47.73 58.83 +0.37
@@ C.1 General Overview
Magpie [Xu et al., 2025] is a self-synthesis framework that derives alignment-oriented annotations from open-weight, instruction-tuned language models without requiring human labels or handcrafted seed prompts. Although the framework is also capable of generating new instruction-response pairs, in this work we rely exclusively on its annotation capabilities.
@@
A key feature of Magpie is its modularity: any instruction-tuned LLM can, in principle, serve as the judge model. In its default configuration, Magpie relies on Llama-3-8B-Instruct [Grattafiori et al., 2024] for general annotation and Llama-Guard 2 [MetaAI, 2024] for safety assessment. In our pipeline, we use Llama-3.3-70B-Instruct [Grattafiori et al., 2024] as the judge model given its demonstrated reliability, particularly for quality and difficulty annotations, as assessed by Djuhera et al. [2025]. In addition, we also incorporate the error-tolerant JSON parsing extensions to Magpie from Djuhera et al. [2025], which help handle formatting inconsistencies and free-form outputs during the annotation process.
@@ C.2 Preference Reward Assessment
We generate preference rewards by adapting Magpie to evaluate the response quality of each completion. To this end, we rely on an independent, specialized reward model that has been fine-tuned on multiple preference datasets. Specifically, we use FsfairX, a Llama-3-8B-Instruct-based reward model introduced by Dong et al. [2024], which was trained using the Bradley-Terry formulation on a diverse set of high-quality preference datasets, including HH-RLHF [Anthropic, 2024], SHP [Ethayarajh et al., 2025], and UltraFeedback Cui et al. [2024]. This reward model has demonstrated strong alignment with human preference judgments, making it a reliable signal for quantifying how well a model completion satisfies the corresponding instruction.
@@
While we acknowledge that other reward models can be used as alternatives, such as UltraRM-13B [Cui et al., 2024], Snorkel-Mistral-PairRM-DPO [Tran et al., 2023], ArmoRM [Wang et al., 2024a], Skywork-Reward [Liu et al., 2024], or more generally, pairwise preference models trained on high-quality preference datasets, Dong et al. [2024] demonstrated through various experiments that their Llama-based FsfairX model performs comparably or more reliably than these alternatives and aligns closely with human judgments, thereby justifying our choice. Nevertheless, we provide a side-by-side comparison to the Skywork reward model for TuluDPO and ORPO datasets in Fig. 6(a) and Fig. 6(b), respectively, showing overall similar distributions for the difference in preference rewards.
In particular, the similar shape of the distributions indicates that preference reward assessment follows the same general trend and does not fundamentally alter the evaluation outcome. At the same time, we observe that the FsfairX model produces fewer negative outliers and more stable margins, making it particularly well-suited for providing reliable training signals without over-amplifying differences. In contrast, the Skywork reward model exhibits both stronger positive separations and more negative outliers, suggesting higher variance. For our purposes, FsfairX provides the consistency and robustness required for a balanced assessment of preference rewards, with its margins further validated against human feedback in several experiments [Dong et al., 2024].